Circular No. 8197

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hv IN NGC 1201 Further to IAUC 8195, B. Beutler and W. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 12.5) on an unfiltered KAIT image taken on Sept. 9.5 UT. The new object was confirmed at mag ∼ 13.0 on an earlier KAIT image taken on Sept. 1.5. SN 2003hv is located at α = 3^{h}04^{m}09^{s}.32, δ = −26^{o}05′07′′.5 (equinox 2000.0), which is 17′′.2 east and 56′′.7 south of the nucleus of NGC 1201. A KAIT image taken on Jan. 28.2 showed nothing at this position (limiting mag ∼ 18.0).

SUPERNOVA 2002lt AND GRB 021211 M. Della Valle, Istituto Nazionale di Astrofisica (INAF), Firenze; D.

Malesani, International School for Advanced Studies, Trieste; S. Benetti, INAF, Padova; and V. Testa and L. Stella, INAF, Rome, on behalf of a larger collaboration (see Della Valle et al. 2003, A.Ap. 406, L33) report the discovery of an apparent supernova associated with GRB 021211 (Crew et al. 2002, GCN 1734) on four R images obtained on 2003 Jan. 9.7 (R = 24.48 ± 0.18, supernova + host galaxy), 15.33 (R = 25.07 ± 0.15), Feb. 28.02 (R = 25.13 ± 0.12), and Mar. 9.01 UT (R = 25.35 ± 0.17), with the Very Large Telescope (VLT; + FORS2) at the European Southern Observatory. SN 2002lt is coincident with the position of the optical afterglow reported by Fox et al. (2003, Ap.J. 586, L5): α = 8^{h}08^{m}59^{s}.86, δ = +6^{o}43′37′′.5 (equinox 2000.0). Inspection of a 4-hr integrated VLT spectrum obtained on Jan. 8.27 (27 days after the GRB outburst) confirms the redshift of the parent galaxy reported by Vreeswijk et al. (2002, GCN 1785) at z = 1.006 and finds broad, low-amplitude undulations blueward and redward of a broad absorption (FWHM ∼ 15.0 nm), the minimum of which is measured at 377.0 nm in the rest frame of the GRB (the blue wing extends up to 365.0 nm). A comparison with spectroscopic templates of supernovae (arranged by different spectroscopic types) strongly supports the identification of the broad absorption with Ca II H + K, blueshifted by ∼ 15000 km/s. The most convincing resemblance is found with the spectra of the ‘prototypical’ type-Ic supernova 1994I (Filippenko et al. 1995, Ap.J. 450, L11) obtained between 0 and 40 days after maximum light.

SATELLITES OF URANUS Corrigendum. On IAUC 8194, line 7, for groundbased read earth-

orbiting

2003 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green